OpenStudy (mathbruh):

NEED HELP MAY GET KICKED OUT OF MATH IF I DONT HAVE ALL DONE PLEASE LAST UESTION FAN MEDAL AND TESTIMONAL

OpenStudy (mathbruh):

Create your own factorable polynomial with a GCF. Rewrite that polynomial in two other equivalent forms. Explain how each form was created.

OpenStudy (alekos):

well you could start with y = (x-1)(x-2)(x-3) a cubic with three factors just expand to get another form

OpenStudy (mathbruh):

I dont know how to do this :(

OpenStudy (mathbruh):

Could you please help me?

OpenStudy (alekos):

(x-1)(x-2)(x-3) = (x^2-3x+2)(x-3) Now keep going with the expansion to create a cubic or 3rd order polynomial

OpenStudy (mathmate):

Not sure if I understand the question: "Create your own factorable polynomial with a GCF." A GCF exists between two different expressions or polynomials, not for a single polynomial.